# Defending Against Unforeseen Failure Modes with Latent Adversarial Training

### v4

@@ Abstract
Despite extensive diagnostics and debugging by developers, AI systems sometimes exhibit harmful unintended behaviors. Finding and fixing these is challenging because the attack surface is so large – it is not tractable to exhaustively search for inputs that may elicit harmful behaviors. Red-teaming and adversarial training (AT) are commonly used to improve robustness, however, they empirically struggle to fix failure modes that differ from the attacks used during training. In this work, we utilize latent adversarial training (LAT) to defend against vulnerabilities without leveraging knowledge of what they are or using inputs that elicit them. LAT makes use of the compressed, abstract, and structured latent representations of concepts that the network actually uses for prediction. Here, we use it to defend against failure modes without examples that elicit them. Specifically, we use LAT to remove trojans and defend against held-out classes of adversarial attacks. We show in image classification, text classification, and text generation tasks that LAT usually improves both robustness to novel attacks and performance on clean data relative to AT. This suggests that LAT can be a promising tool for defending against failure modes that are not explicitly identified by developers. [^1] Code is available at https://github.com/thestephencasper/latent_adversarial_training.
**footnotetext: Equal contribution.
@@ 1 Introduction
Ensuring that AI systems will be trustworthy, even in the face of anomalous and adversarial inputs, has been a major focus of research in the past decade (Szegedy et al., 2013; Goodfellow et al., 2014; Zhao et al., 2022; Anwar et al., 2024; Yohsua et al., 2024), and it has been incorporated into risk management frameworks for AI governance (NIST, 2023; DSIT, 2023; EU, 2021; NISSTC, 2023). Developers commonly use test sets, red-teaming, and attack methods to identify vulnerabilities followed by adversarial training (AT) to fix them. This is valuable, but sometimes fails to address problems. There are often systematic differences between the failure modes that developers search for (e.g., $L_{p}$-norm attacks) and ones that models can exhibit post-deployment (e.g. (Hendrycks et al., 2021a, c)). Many real-world vulnerabilities can evade detection such as trojans (Hubinger et al., 2024; Carlini et al., 2022), jailbreaks (Liu et al., 2023; Wei et al., 2023; Zou et al., 2023b; Shah et al., 2023) novel attacks (Brown et al., 2018; Shayegani et al., 2023; Geiping et al., 2024), or black swans (Kolt, 2023; Hendrycks et al., 2021b). In Figure 1a, we illustrate the gap between failure modes developers identify, and the “unforeseen” ones they do not.
@@
Standard attack and red-teaming techniques require searching a model’s input space for examples that elicit failures. This is challenging – the input space is massive, and it is easy for some failure modes to go unnoticed (Goel et al., 2024). In this paper, we use latent adversarial training (LAT) (Sankaranarayanan et al., 2018) as an additional way to defend against failures without requiring examples that trigger them. In contrast to AT which uses attacks in the input space, LAT uses attacks in the latent representations. Figure 1c illustrates this distinction. LAT has previously been used as a computationally efficient way to improve robustness to conventional $L_{p}$ norm attacks while limiting harm to performance on clean data (e.g., (Sankaranarayanan et al., 2018; Singh et al., 2019; Park and Lee, 2021; Kitada and Iyatomi, 2023), see Section 2 for a full overview). However, here, we specifically study its ability to reduce unforeseen risks.
@@
Our motivation for LAT is based on the vastness of a model’s input space compared to the manifold of task-relevant features within it. Across the latents, a model gradually develops more compressed, abstract, and structured representations of the concepts it uses to process information. [^2] Since the mappings from layer to layer are not typically surjective, not all latent states are inducible by real inputs. This relaxation of the problem is what gives LAT the potential to compensate for the gap between failure modes that are identified by developers and ones that are not (see Figure 1a). This makes it possible for latent space attacks to activate neural circuitry that elicits failures (Figure 1b) without requiring inputs that trigger them (Fort, 2023). Thus, we hypothesize that even if it is difficult to find a problem with a model using input space attacks, it may be comparatively easier with latent space ones. We make three contributions:
Figure 1: (a) We study latent adversarial training (LAT) as a method to reduce risks from failure modes that are not identified by developers pre-deployment. (b) Our motivation for LAT is based on how models develop more compressed, abstract, and structured representations across their latents. We hypothesize that many failures that are difficult to elicit from the input space may be easier to elicit from the latent space. (c) In experiments, we compare AT with LAT. The key difference is that AT applies adversarial perturbations to the input, and LAT applies adversarial perturbations to a hidden layer. (d) In each of our experiments, we compare methods based on (1) their performance on clean evaluation data, (2) their robustness to novel classes of adversarial examples not encountered during training, and (3) their robustness to trojans implanted during pretraining. We find that LAT in the optimal layer usually performs better than AT under all three evaluations.
@@
    We show in vision, language understanding, and language generation tasks that LAT can improve robustness against failure modes without any examples that elicit them. Specifically, we use it to defend against trojans and novel classes of attacks. We find that LAT in the optimal layer usually Pareto dominates AT with respect to both clean and robust performance.
@@
    We demonstrate cautionary instances in which robustness techniques sometimes harm robustness to novel failure modes. We show an instance in which $L_{p}$-norm AT in vision models makes a network more susceptible to novel attacks. Also similar to recent findings from Hubinger et al. (2024), we show instances in which, under certain configurations, AT and LAT (without using examples containing a trojan trigger) can make a trojaned LLM more susceptible to its trojan.
@@ 2 Related Work
  • Trojans (also known as backdoors) can be triggered by arbitrary features (Chen et al., 2017; Wu et al., 2022; Carlini et al., 2023).
@@
  • Black swans refer to anomalies which avoid detection due to their rarity (Kolt, 2023; Hendrycks et al., 2021b).
@@
Latent space attacks in vision models: Several works have experimented with latent space attacks and LAT at small scales (Singh et al., 2019; Park and Lee, 2021; Qian et al., 2021; Zhang et al., 2023) while (Sankaranarayanan et al., 2018) did so at the ImageNet scale. Several of these works have found that LAT can improve robustness to $L_{p}$-norm input space attacks and generalization on clean data (Sankaranarayanan et al., 2018; Singh et al., 2019). However, in contrast to any of the above, we use LAT to increase robustness to more novel failure modes in the form of trojans and non-$L_{p}$-norm attacks.
@@
Latent space attacks in language models: In language models, it is not possible to directly use gradient-based methods to generate adversarial attacks because tokenization is not differentiable. However, several works have attacked word embeddings (which can be viewed as the first latent state in the network) and trained on these perturbations to improve robustness or generalization (Jiang et al., 2019; Zhu et al., 2019; Liu et al., 2020; He et al., 2020; Kuang and Bharti, ; Li and Qiu, 2021; Sae-Lim and Phoomvuthisarn, 2022; Pan et al., 2022; Schwinn et al., 2023; Geisler et al., 2024; Schwinn et al., 2024; Xhonneux et al., 2024). Here, we use embedding space adversarial training as a baseline. Deeper in the latent space, Fort (2023) demonstrated that language models can be very sensitive to latent perturbations. Others have shown that LLMs can have their high-level behaviors altered by perturbations to their latent staftes found through probing or causal interventions (Turner et al., 2023; Li et al., 2023; Zou et al., 2023a; Wang and Shu, 2023; Rimsky et al., 2023; Jorgensen et al., 2023; Lu and Rimsky, 2024; von Rütte et al., 2024). However, to the best of our knowledge, these types of perturbations have not yet been used for LAT. The most similar work to ours is Kitada and Iyatomi (2023), who perform LAT on attention representations to improve generalization in small BERT-scale models. In contrast, however, we use models up to 7B parameters in size, use more general attacks by targeting the residual stream of transformers, test both classifiers and generators, and evaluate methods under trojans and novel attacks.
@@ 3 Method
Threat model: The threat that we consider is not an attacker that has access to the latent space. Although we train under latent perturbations, our ultimate goal is not to make the trained model resistant to these. Instead, our goal is to make models robust to distribution shifts between development and deployment that are not precisely known beforehand (e.g., trojans, jailbreaks, novel attacks, and black swans).
@@
Note that this setup involves ‘untargeted’ attacks in which the adversary maximizes the target model’s loss. ‘Targeted’ attacks in which the adversary elicits a specific target output are possible, but we leave this to future work.
For all attacks, we clip the perturbed activations by the min and max of the unperturbed activations across all neurons in the current batch to reduce the risk of attacks moving activations to an irrelevant part of the latent space.
Latent space distance metric: Typically, AT constrains the perturbation according to a constraint defined by a simple distance metric such as an $L_{p}$-norm. This is reasonable for AT because all input components (e.g., pixels) have comparable activation distributions. However, this is not guaranteed for latent representations – each neuron may have a different distribution of activations. As a result, we experiment with using a normalized distance metric to constrain latent perturbations. However, we find no clear differences in results between using standard and normalized distance metrics. In Section 4, we pool results using standard and normalized distance metrics together, but in the Appendix, we present each side-by-side.
@@ 4 Experiments
Three tasks: image classification, text classification, and text generation. We experiment with three different tasks: image classification on ImageNet (Russakovsky et al., 2014), text classification on the Redwood injurious text dataset (Ziegler et al., 2022), and text generation on the Anthropic Helpful-Harmless-RLHF (Bai et al., 2022) and PKU BeaverTails (Ji et al., 2024) data distributions.
@@
Three methods: AT, LAT, and RLP. In each experiment, we compare 3 methods: AT, LAT, and training under random latent perturbations (RLP). We use these random latent perturbations with the same norm as LAT perturbations as a non-adversarial contrast to LAT. We select the latent layer to perturb by sweeping across layers for high clean and robust performance. We converged to the heuristics of perturbing the first post-convolutional layer in CNNs and a relatively early layer in transformers. We produce all attacks using projected gradient descent (Madry et al., 2017).
@@
Three measures: clean performance, adversarial robustness, and trojan forgetting. For each experiment, we first fine-tune the model using poisoned data to implant trojans. Second, we fine-tune further on clean data while applying RLP, AT, and LAT. We report results across this second fine-tuning stage. For each task, we evaluate methods based on (1) performance on clean data, (2) robustness to novel classes of input-space adversarial attacks, and (3) robustness to trojans we implanted. We illustrate this in Figure 1d.
@@
Our goal: Expanding the Pareto frontier. Because in different applications, practitioners may prefer different tradeoffs between clean and robust performance, we focus on the Pareto frontier between clean performance and each type of robustness. In each experiment, we perform multiple runs of RLP, AT, and LAT with varying perturbation bounds ($\epsilon$) and evaluate at multiple training checkpoints for each. This allows us to construct a scatterplot of different tradeoffs between clean and robust performance. Overall, we find that LAT in the optimal layer usually Pareto dominates RLP and AT with respect to both clean and robust performance.
@@
Figure 2: Image classification: Latent adversarial training Pareto-dominates adversarial training and random latent perturbations with respect to clean and robust accuracy. Points further up and to the right are better. To show which checkpoints came from the same training run, we connect sets by thin lines. (a) Clean accuracy change compared to robust accuracy change on novel classes of attacks from Kaufmann et al. (2019). (b) Clean accuracy change compared to robust accuracy change under previously implanted trojans.
@@ 4.1 Image Classification
We used a ResNet-50 from He et al. (2016) and fine-tuned it on a version of the ImageNet (Russakovsky et al., 2014) training set that was poisoned as in Casper et al. (2023) to implant 8 trojans: four with a $64\times 64$ pixel patch trigger, and four with a natural feature trigger which took the form of a pre-existing feature in certain images. In both cases, the trojan response was a randomly selected label different from the ground truth. We then fine-tuned the model on clean ImageNet training data for one epoch using RLP, AT, and LAT. When performing LAT, we attacked the first layer activations after the residual blocks (pre-avgpooling). Finally, we evaluated the resulting models on (1) the clean ImageNet test set, (2) the ImageNet test set images, each attacked using one of the 18 held-out attack methods from Kaufmann et al. (2019), [^3] We used 18 of the 20 attacks from Kaufmann et al. (2019) excluding FGSM and PGD to avoid contaminating the test set with $L_{p}$-norm input-space attacks. and (3) the ImageNet test set images, each attacked with one of the 8 trojan triggers.
We plot (1) vs (2) and (1) vs (3) in Figure 2. All Pareto frontiers are entirely filled with model checkpoints from LAT. Compared to AT, LAT methods result in 68%, 68%, and 1.5% greater improvements to the area under the Pareto curve for attack robustness (a) patch trojan removal (b.1) and natural feature trojan removal (b.2) respectively. We also find an example of how $L_{p}$-norm AT can be harmful to robustness against novel attacks (see Figure 2a and Figure 5a). AT using $L_{p}$-norm attacks caused the model to be more susceptible to other attacks from Kaufmann et al. (2019). [^4] Kaufmann et al. (2019) found that $L_{p}$ norm AT could be used to improve robustness to their attacks. Our results do not conflict with this. We find in Figure 2 that some checkpoints and some runs of AT did improve robustness, but we nonetheless found that most checkpoints from most AT runs were less robust than before AT.

### v6

@@ Abstract
Despite extensive diagnostics and debugging by developers, AI systems sometimes exhibit harmful unintended behaviors. Finding and fixing these is challenging because the attack surface is so large – it is not tractable to exhaustively search for inputs that may elicit harmful behaviors. Red-teaming and adversarial training (AT) are commonly used to improve robustness, however, they empirically struggle to fix failure modes that differ from the attacks used during training. In this work, we utilize latent adversarial training (LAT) to defend against vulnerabilities without leveraging knowledge of what they are or using inputs that elicit them. LAT makes use of the compressed, abstract, and structured latent representations of concepts that the network actually uses for prediction. Here, we use it to defend against failure modes without examples that elicit them. Specifically, we use LAT to remove backdoors and defend against held-out classes of adversarial attacks. We show in image classification, text classification, and text generation tasks that LAT usually improves both robustness to novel attacks and performance on clean data relative to AT. This suggests that LAT can be a promising tool for defending against failure modes that are not explicitly identified by developers. [^1] Code is available at https://github.com/thestephencasper/latent_adversarial_training. See also https://github.com/aengusl/latent-adversarial-training.
Figure 1: (a) We study latent adversarial training (LAT) as a method to reduce risks from failure modes that are not identified by developers pre-deployment. (b) Our motivation for LAT is based on how models develop more compressed, abstract, and structured representations across their latents. We hypothesize that many failures that are difficult to elicit from the input space may be easier to elicit from the latent space. (c) In experiments, we compare AT with LAT. The key difference is that AT applies adversarial perturbations to the input, and LAT applies adversarial perturbations to a hidden layer. (d) In each of our experiments, we compare methods based on (1) their performance on clean evaluation data, (2) their robustness to novel classes of adversarial examples not encountered during training, and (3) their robustness to backdoors implanted during pretraining. We find that LAT in the optimal layer usually performs better than AT under all three evaluations.
@@ 1 Introduction
Ensuring that AI systems will be trustworthy, even in the face of anomalous and adversarial inputs, has been a major focus of research in the past decade (Szegedy et al., 2013; Goodfellow et al., 2014; Zhao et al., 2022; Anwar et al., 2024; Yohsua et al., 2024), and it has been incorporated into risk management frameworks for AI governance (NIST, 2023; DSIT, 2023; EU, 2024; NISSTC, 2023). Developers commonly use test sets, red-teaming, and attack methods to identify vulnerabilities followed by adversarial training (AT) to fix them. This is valuable, but sometimes fails to address problems. There are often systematic differences between the failure modes that developers search for (e.g., $L_{p}$-norm attacks) and ones that models can exhibit post-deployment (e.g. (Hendrycks et al., 2021a; c)). Many real-world vulnerabilities can evade detection such as backdoors (Hubinger et al., 2024; Carlini et al., 2022), jailbreaks (Liu et al., 2023; Wei et al., 2023; Zou et al., 2023b; Shah et al., 2023) novel attacks (Brown et al., 2018; Shayegani et al., 2023; Geiping et al., 2024), or black swans (Kolt, 2023; Hendrycks et al., 2021b). In Figure˜1a, we illustrate the gap between failure modes developers identify and the “unforeseen” ones they do not.
@@
Standard attack and red-teaming techniques require searching a model’s input space for examples that elicit failures. This is challenging – the input space is massive, and it is easy for some failure modes to go unnoticed (Goel et al., 2024). In this paper, we use latent adversarial training (LAT) (Sankaranarayanan et al., 2018) as an additional way to defend against failures without requiring examples that trigger them. In contrast to AT which uses attacks in the input space, LAT uses the same attack method on the latent representations. Figure˜1c illustrates this distinction. LAT has previously been used as a computationally efficient way to improve robustness to conventional $L_{p}$ norm attacks while limiting harm to performance on clean data (e.g., (Sankaranarayanan et al., 2018; Singh et al., 2019; Park & Lee, 2021; Kitada & Iyatomi, 2023), see Section˜2 for a full overview). However, here, we specifically study its ability to reduce novel risks. [^2] Since LAT uses the same PGD attack method as AT, it is a version of AT and is able to play the same role as AT in training pipelines. LAT is, therefore, complementary to existing latent space methods such as Minh & Tuan (2022); Yang et al. (2024); Zhou et al. (2021); Moon et al. (2023); Rosati et al. (2024); Wang et al. (2021).
@@
Our motivation for LAT is based on the vastness of a model’s input space compared to the manifold of task-relevant features within it. Across the latents, a model gradually develops more compressed, abstract, and structured representations of the concepts it uses to process information (Wang et al., 2023).333 Multiple inputs may map to similar latent states. Therefore, finding and defending against implicit failure modes in the latent neighborhood of a given sample may correspond to distant or unknown inputs, bridging the gap between failure modes that are identified by developers and ones that are not (see Figure 1a). See also prior non-archival discussions of this principle from Christiano (2019); Hubinger (2019); Jermyn (2022). This makes it possible for latent-space attacks to activate neural circuitry that elicits failures (Figure˜1b) without requiring inputs that trigger them (Fort, 2023). Thus, we hypothesize that even if it is difficult to find a weakness with a model using input-space attacks, it may be comparatively easier with latent-space ones. Accordingly, this paper seeks to study whether LAT can confer more generalizable forms of robustness than input-space AT. We make three contributions:
@@
    We show in vision, language understanding, and language generation tasks that LAT can improve robustness against failure modes without any examples that elicit them. Specifically, we use it to defend against backdoors and novel classes of attacks. We find that LAT in the optimal layer usually Pareto-dominates AT with respect to both clean and robust performance.
@@
    We demonstrate cautionary instances in which robustness techniques sometimes harm robustness to novel failure modes. We show an instance in which $L_{p}$-norm AT in vision models makes a network more susceptible to novel attacks. Also, similar to recent findings from Hubinger et al. (2024), we show instances in which, under certain configurations, AT and LAT (without using examples containing a backdoor trigger) can make a backdoored LLM more susceptible to its backdoor.
@@ 2 Related Work
  • Backdoors (also known as trojans) can be triggered by arbitrary features (Chen et al., 2017; Wu et al., 2022; Carlini et al., 2023).
@@
  • Black swans refer to harmful anomalies which avoid detection due to their rarity (Kolt, 2023; Hendrycks et al., 2021b).
@@
[…]
@@
Latent-space attacks in language models: In language models, it is not possible to directly use gradient-based methods to generate adversarial attacks because tokenization is not differentiable. However, several works have attacked word embeddings (which can be viewed as the first latent state in the network) and trained on these perturbations to improve robustness or generalization (Jiang et al., 2019; Zhu et al., 2019; Liu et al., 2020; He et al., 2020; Kuang & Bharti, ; Li & Qiu, 2021; Sae-Lim & Phoomvuthisarn, 2022; Pan et al., 2022; Schwinn et al., 2023; Geisler et al., 2024; Schwinn et al., 2024; Xhonneux et al., 2024). Here, we use embedding-space adversarial training as a baseline. Deeper in the latent space, Fort (2023) demonstrated that language models can be very sensitive to latent perturbations. Others have shown that LLMs can have their high-level behaviors altered by perturbations to their latent states found through probing or causal interventions (Turner et al., 2023; Li et al., 2023; Zou et al., 2023a; Wang & Shu, 2023; Rimsky et al., 2023; Jorgensen et al., 2023; Lu & Rimsky, 2024; von Rütte et al., 2024). However, to the best of our knowledge, these types of perturbations have not yet been used for LAT. Furthermore, a series of recent papers have explored fine-tuning models under perturbations to weights or activations to make them more resilient to unwanted downstream fine-tuning (Henderson et al., 2023; Deng et al., 2024; Huang et al., 2024b; a; Tamirisa et al., 2024; Rosati et al., 2024; Huang et al., 2024c). Finally, our work is the most similar to Kitada & Iyatomi (2023), who perform LAT on attention representations to improve generalization in small BERT-scale models, and concurrent work from (Huang et al., 2024d), who use LAT to make models more resistant to unwanted fine-tuning in LLMs. In contrast, however, we are the first to study LAT’s ability in LLMs to remove existing harmful behaviors. We also evaluate methods under both backdoors and novel attacks.
@@ 3 Method
Threat model: The threat that we consider is not an attacker that has access to the latent space. Although we train under latent perturbations, our ultimate goal is not to make the trained model resistant to these. Instead, our goal is to make models robust to distribution shifts between development and deployment that are not precisely known beforehand (e.g., backdoors, jailbreaks, novel attacks, and black swans).
@@
Note that this setup involves ‘untargeted’ attacks in which the adversary maximizes the target model’s loss. ‘Targeted’ attacks in which the adversary elicits a specific target output are possible, but we leave this to future work. We present the full LAT algorithm in Appendix˜B.
Latent space distance metric: Typically, AT constrains the perturbation according to a constraint defined by a simple distance metric such as an $L_{p}$-norm. This is reasonable for AT because all input components (e.g., pixels) have comparable activation distributions. However, this is not guaranteed for latent representations – each neuron may have a different distribution of activations. As a result, we experiment with using a normalized distance metric to constrain latent perturbations. However, we find no clear differences in results between using standard and normalized distance metrics. In Section˜4, we pool results using standard and normalized distance metrics together, but in Appendix˜E, we present each side-by-side.
@@ 4 Experiments
Three tasks: image classification, text classification, and text generation. We experiment with three different tasks: image classification on ImageNet (Russakovsky et al., 2014), text classification on the Redwood injurious text dataset (Ziegler et al., 2022), and text generation on the Anthropic Helpful-Harmless-RLHF (Bai et al., 2022) and PKU BeaverTails (Ji et al., 2023) data distributions.
@@
Three methods: AT, LAT, and RLP. In each experiment, we compare three methods: AT, LAT, and training under random latent perturbations (RLP). We use these random latent perturbations with the same norm constraint as LAT perturbations as a non-adversarial contrast to LAT. We select the latent layer to perturb by sweeping across layers for high clean and robust performance. We converged to the heuristics of perturbing the first post-convolutional layer in CNNs and a relatively early layer in transformers. We produce all attacks using projected gradient descent (PGD) (Madry et al., 2017). [^4] On efficiency: AT is the most computationally expensive, followed by LAT and then RLP. AT requires $T_{\delta}$ forward and backward passes through the model to develop an attack that is optimized for $T_{\delta}$ steps, additional to the forward and backward pass of the regular training step. LAT is somewhat cheaper. While it still requires $T_{\delta}$ forward and backward passes, the passes start from and end at the target layer. The efficiency gains will depend on what target layer is used. Finally, RLP is the most computationally cheap, requiring no adversarial optimization.
@@
Three measures: clean performance, adversarial robustness, and backdoor removal. For each experiment, we first fine-tune the model using poisoned data to implant backdoors. Second, we fine-tune further on clean data while applying RLP, AT, and LAT. We report results across this second fine-tuning stage. For each task, we evaluate methods based on (1) performance on clean data, (2) robustness to novel classes of input-space adversarial attacks, and (3) robustness to backdoors implanted through data poisoning. [^5] We use straightforward data poisoning with conspicuous examples. However, other, more subtle techniques for implanting backdoors exist as well (Guo et al., 2022). We illustrate this in Figure˜1d. While not unforeseen to us, we use held-out attacks and backdoors as proxies for “unforeseen” failure modes because they are fully held out from the adversarial training process. This methodological use of backdoors is similar to prior work (Hubinger et al., 2024; Hofstätter et al., 2025). We do not compare LAT to backdoor-specific defense methods (Zhao et al., 2024) because our goal is to study LAT’s ability to defend against unforeseen vulnerabilities in general.
@@
Our goal: Expanding the Pareto frontier. Because in different applications, practitioners may prefer different tradeoffs between clean and robust performance, we focus on the Pareto frontier between clean performance and each type of robustness. In each experiment, we perform multiple runs of RLP, AT, and LAT with varying perturbation constraints ($\epsilon$) and evaluate at multiple training checkpoints for each. [^6] We use the same effort to tune the hyperparameters that overlap between AT and LAT. Tuning LAT is the same as tuning AT except for the additional hyperparameter of what layer to attack. This allows us to construct a scatterplot of different tradeoffs between clean and robust performance. Overall, we find that LAT in the optimal layer usually Pareto-dominates RLP and AT with respect to both clean and robust performance.
@@
Figure 2: Image classification: Latent adversarial training Pareto-dominates adversarial training and random latent perturbations with respect to clean and robust accuracy. Points further up and to the right are better. To show which checkpoints came from the same training run, we connect sets by thin lines. (a) Clean accuracy compared to robust accuracy on novel classes of attacks from Kaufmann et al. (2019). (b) Clean accuracy compared to robust accuracy under previously implanted backdoors.
@@ 4.1 Image Classification
We used a ResNet-50 from He et al. (2016) and fine-tuned it on a version of the ImageNet (Russakovsky et al., 2014) training set that was poisoned as in Casper et al. (2023) to implant 8 backdoors: four with a “patch” trigger and four with a “natural feature” trigger Casper et al. (2023). All 8 backdoors had a randomly-selected target class. We then fine-tuned the model on clean ImageNet training data for one epoch using RLP, AT, and LAT. When performing LAT, we attacked the first layer activations after the residual blocks (pre-avgpooling). We evaluated the resulting models on (1) the clean ImageNet test set, (2) the ImageNet test set images, each attacked using one of the 18 held-out attack methods from Kaufmann et al. (2019), [^7] We used 18 of the 20 attacks from Kaufmann et al. (2019) excluding FGSM and PGD to avoid contaminating the test set with $L_{p}$-norm input-space attacks. and (3) the ImageNet test set images, each attacked with one of the 8 backdoor triggers.
We plot (1) vs (2) and (1) vs (3) in Figure˜2. All Pareto frontiers are entirely filled with model checkpoints from LAT. Compared to AT, LAT methods result in 68%, 68%, and 1.5% greater improvements to the area under the Pareto curve for novel attack robustness (a) patch backdoor removal (b.1) and natural feature backdoor removal (b.2), respectively. We also find an example of how $L_{p}$-norm AT can be harmful to robustness against novel attacks (see Figure˜2a and Figure˜5a). [^8] This may be related to findings from Ilyas et al. (2019) which could explain cases in which AT substantially harms both clean and robust accuracy. It is possible that $L_{p}$-norm AT harms the model’s ability to pick up on useful $L_{p}$-norm features that are not manipulated by the held-out attacks. AT using $L_{p}$-norm attacks caused the model to be more susceptible to other attacks from Kaufmann et al. (2019). [^9] Kaufmann et al. (2019) found that $L_{p}$ norm AT could be used to improve robustness to their attacks. Our results do not conflict with this. We find in Figure 2 that some checkpoints and some runs of AT did improve robustness, but we nonetheless found that most checkpoints from most AT runs were less robust than before AT.
Figure 3: Text classification: Latent adversarial training improves over embedding-space adversarial training across much of the Pareto frontier. Points further up and to the right are better. We connect evaluation checkpoints from the same run by thin lines. (a) Clean ROC-AUC compared to robust ROC-AUC on unseen manually-generated attacks from Ziegler et al. (2022). (b) Clean ROC-AUC compared to backdoor ROC-AUC. Overall, LAT does not always outperform AT and RLP but does so in “elbow” cases that most evenly balance clean and robust performance.
Figure 4: Text generation: (a, c) Latent adversarial training matches or improves on embedding-space adversarial training for forgetting undesirable text. (b, d) However, for removing backdoors, the dataset used affected which method performs better. Points further up and to the right are better. We connect evaluation checkpoints from the same run by thin lines. (a, c) Loss on preferred/harmless examples compared against loss on rejected/harmful data. For Anthropic-HH, the model’s performance on the preferred versus rejected distributions is highly correlated. For the BeaverTails dataset, LAT Pareto-dominates AT. (b, d) Loss on preferred/harmless data compared against loss on backdoors. Despite the same backdoors being used in each case, LAT dominates on Anthropic-HH while AT dominates on BeaverTails. Backdoor removal on BeaverTails is the only experiment in which we find AT to broadly outperform LAT.